At 234 degrees Celsius, the K eq for the reaction is 4.93x 1 ^1 1A+5B leftrightarrow 4AB. What is kp?

At 234 °C, the Key for the reaction is 4.93 · 101: -1 A+5B H 4 AB What is Kp? Show your work:


At 234 °C, the Kp for the reaction is 4.93 . 10: 1A + 5B # 4AB What is Keq?
Answer #1

Answer:

1. K​​​​​​P = 2.77 × 10-6

2. K​​​​​​eq = 8.76 × 108

Explanation:

To solve these questions we need the following formula,

K​​​​​​P = K​​​​​eq (RT)\Deltan

Where,

R = universal gas constant

T = temperature

\Deltan = (number of moles of products - number of moles of reactants)

1.

1 A +5B 4 AB No. of moles of product = 4 No. of moles of reactant - 1+5=6 So, An= 4-6=-2 Keq = 4.93×10² T= 234°C = 507.15 K R

2.

For the second question, we will use the same equation.

1A +5B 4AB

Here,

An = -2 Kp = 4.938 10² = 507.15K R = 8.314 J.K. mal put these values in the equation for Kp 4.934 18 ² = keq (8.3144 507.15